What is the pH of a solution of HCl with a concentration of 1.2 X 10
-3 M?
1
Expert's answer
2019-07-01T03:54:51-0400

Solution.

Hydrochloric acid is a strong acid, therefore, to find the acidity (pH), the following formula is necessary:



"pH = -lg(H+)""pH = -lg(1.2 \times 10^{-3})"

pH = 2.92

Answer:

pH = 2.92